OpenStudy (calchelp):

IS the sum of (3^n - 2^2)/ 2^(2n) convergent and what is the number. pls show steps.

OpenStudy (anonymous):

it converges as : =(3^n-2^2)/4^n =3^n/4^n-4/4^n =(3/4)^n-4^(1-n) so |3/4|=.75<1 so it converges to zero when it is raised to infinity and 4^(1-n) when raised to infinity it converges to zero so the equation converges to zero

OpenStudy (calchelp):

thanks can you also find the interval of convergence.